OpenStudy (anonymous):

2 3+a 2 -- - --- + --------- = a+3 2a+6 a²+6a+9a²

OpenStudy (anonymous):

first you need to make a common denominator to add all these fractions. a²+6a+9a²=(a+3)^2 2a+6=2(a+3) so we can see that the least common denominator will be 2* (a+3)^2 Now you have to expand all there will be (2*2*(a+3) -(3+a)(a+3) +4)/(2(a+3)^2 4a+12-a^2-6a-9+4/(2(a+3)^2 (-a^2-2a+7)/(2(a+3)^2)

OpenStudy (anonymous):

are you sure this is correct. I do not have that answer

OpenStudy (anonymous):

no it looks ugly, probably I did some mistake, I am too tired. I really should go to bed

OpenStudy (anonymous):

hope that the idea behind it is clear

OpenStudy (anonymous):

-2 3a+7 -- - --- --------- 2(a+3)(a+3) 2a²+12a+18 -1 9a+25 ---------- --------- 2(a²+6a+9) 2(a+3)² these are my choice,which is the correct one?

OpenStudy (anonymous):

have a goodnight